OpenStudy (anonymous):

what is three-quarters of seven ninths,expressed as a fraction in its simplest form

OpenStudy (alekos):

fractional arithmetic 3/4 x 7/9 = 21/36 = 7/12

OpenStudy (anonymous):

A quarter is : 1/4, so three quarters will be : 3/4.. So; 3/4 of 7/9 = (3*7)/(4*9)

OpenStudy (anonymous):

\[\frac{3}{4} \times \frac{7}{9} = \frac{1}{4} \times \frac{7}{3} = \color{green}{\frac{7}{12}}\]